OpenStudy (anonymous):

Find the 5th term of the sequence in which t1 = 8 and tn = -3tn-1.

OpenStudy (anonymous):

8 = 3n - 1 9 = 3n n = 3\[\sum_{3}^{7} (3n - 1)\] \[a _{5} = 3n - 1 = 21 - 1 = 20\]

OpenStudy (anonymous):

I have one question left: Find the 6th term of the sequence with t1 = -4 and tn = 5tn-1.

OpenStudy (anonymous):

I'm not very sure with this one...sorry.

OpenStudy (anonymous):

oh im sorry i meant this Find the 6th term of the sequence with t1 = -4 and tn = 5^tn-1

OpenStudy (anonymous):

the last one is squared

OpenStudy (anonymous):

So, \[t _{n} = 5^{n - 1}?\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

I think you should ask someoone else because this involves logs and I'm not good at them.